### v2

# FMNV: A Dataset of Media-Published News Videos for Fake News Detection

Abstract
News media, particularly video-based platforms, have become deeply embedded in daily life, concurrently amplifying risks of misinformation dissemination. Consequently, multimodal fake news detection has garnered significant research attention. However, existing datasets predominantly comprise user-generated videos characterized by crude editing and limited public engagement, whereas professionally crafted fake news videos disseminated by media outlets—often politically or virally motivated—pose substantially greater societal harm. To address this gap, we construct FMNV, a novel dataset exclusively composed of news videos published by media organizations. Through empirical analysis of existing datasets and our curated collection, we categorize fake news videos into four distinct types. Building upon this taxonomy, we employ Large Language Models (LLMs) to automatically generate deceptive content by manipulating authentic media-published news videos. Furthermore, we propose FMNVD, a baseline model featuring a dual-stream architecture integrating CLIP and Faster R-CNN for video feature extraction, enhanced by co-attention mechanisms for feature refinement and multimodal aggregation. Comparative experiments demonstrate both the generalization capability of FMNV across multiple baselines and the superior detection efficacy of FMNVD. This work establishes critical benchmarks for detecting high-impact fake news in media ecosystems while advancing methodologies for cross-modal inconsistency analysis. Our dataset is available in https://github.com/DennisIW/FMNV.
Keywords:
Multimodal Fake News Detection News Video Dataset Large Language Model Data Augmentation Deep Learning.
1 Introduction
With the rapid evolution of social media into video-centric content ecosystems, multimodal fake news has emerged as a novel information threat in the digital age. By seamlessly integrating textual narratives, audiovisual materials, and contextualized editing techniques, such content constructs highly deceptive narrative [1, 2, 3, 4, 5], exhibiting exponentially greater propagation efficacy compared to traditional unimodal forms. Recent studies reveal that 73% of disinformation on social platforms in 2023 employed cross-modal coordinated manipulation strategies [6], posing severe challenges to public cognitive security. Multimodal Fake News Detection (MFND), as a pivotal task in cross-modal intelligent computing, aims to achieve precise identification of deceptive content through in-depth exploration of semantic correlations and contradictory patterns across text, visual, auditory, and other heterogeneous data sources [2, 12].
The rise of social media based on video has gradually shifted the focus of multi-modal false news detection to the detection of false news videos. However, the existing video data sets all contain many poorly edited videos posted by individual users, which attract a small amount of attention and are easy for people to distinguish the truth and falseness of news, so there is little social harm. However, because the news media receives more attention, its fake news videos that are driven by political propaganda or traffic are more harmful. To address this challenge, we propose Fake Media News Videos (FMNV), a dataset consisting of 2,393 news videos released by news media enhanced with large language model data. Specifically, we conduct an empirical analysis to categorize fake news videos and give examples. Then, we collect 2,393 news videos published by news media on YouTube and Twitter. Since the workload of human annotation is too large and the accuracy is not high, and most of the videos published by media are real news videos, there is a problem of data imbalance, so we adopt Large Language Models (LLMs) for data enhancement. 1,500 of the news videos are changed to different categories of fake news videos. Each video contains three types of modal information: title, video clip and audio. Fake news videos have the problem of inconsistent information between different modals or consistent information between modals but contrary to common sense.
We also propose Fake Media News Video Detection (FMNVD) as the baseline for FMNV. This model employs a dual-stream architecture integrating CLIP and Faster R-CNN to capture discriminative video features, followed by a co-attention layer for multimodal aggregation. Numerical experiments conducted on FakeTT and FMNV datasets using existing methods and FMNVD demonstrate the effectiveness of our proposed framework.
The main contributions are summarized as follows:
  • $\bullet$
    We conduct an empirical analysis to categorize fake news videos in video datasets and the real world, giving examples of each category.
  • $\bullet$
    We construct FMNV, a novel fake news video detection dataset that contains only news videos published by news media. And we use LLMs data augmentation instead of manual annotation, solving the problem of data imbalance.
  • $\bullet$
    We propose FMNVD, a baseline model that leverages CLIP and Faster R-CNN for feature extraction and employs co-attention for feature aggregation. Comparative experiments validate the effectiveness of FMNV.
2 Related Work
The field of multimodal fake news detection (MFND) has witnessed significant progress through the development of diverse datasets. Early research primarily focused on image-text multimodal scenarios, leading to several influential datasets. FVC-2018 [8] pioneered multimodal verification by incorporating images, text, and metadata from news articles. FakeNewsNet [9] further enriched the landscape with social context features including user profiles and propagation networks. Subsequent innovations introduced specialized datasets: VAVD [10] addressed visual claim verification through image-claim pairs, while LIAR [11] emphasized textual metadata analysis with speaker-specific features. The recently proposed MCFEND [12] advanced cross-modal fusion capabilities through its multi-granularity framework. These datasets collectively established foundational resources for analyzing textual, visual, and social modalities in static content.
However, the emergence of video-centric social platforms has necessitated a paradigm shift toward dynamic multimodal analysis. Researchers have responded by developing video-specific MFND datasets that account for temporal and audiovisual complexities. A notable contribution comes from Qi et al. [3], who constructed FakeSV - a comprehensive Chinese short video dataset containing 5,538 videos from Douyin and Kuaishou platforms. This dataset uniquely categorizes content into 1,827 real videos, 1,827 fake videos, and 1,884 debunked videos across 738 events. Each entry provides rich multimodal features including video content, titles, metadata (upload dates/locations), user comments, and creator profiles. In 2024, Bu and Qi et al. [14], based on FakeSV, proposed FakeTT, an annotated English video dataset with audiovisual text inconsistencies from TikTok. These video datasets emphasize dynamic multimodal interactions, temporal inconsistency detection, and platform-native engagement features, providing critical infrastructure for detecting evolving deception tactics in video-first media environments.
3 The FMNV Dataset
In this section, we construct a comprehensive dataset through a multi-stage process, integrating empirical analysis, real-world data collection, large-model-assisted augmentation, and in-depth dataset evaluation. First, we examine the semantic inconsistency patterns commonly found in fake news videos to guide data annotation. Then, we collect and curate video samples from various sources, followed by data augmentation using large language models and shot segmentation. Finally, we perform quantitative and qualitative analysis, comparing our dataset with existing benchmarks to highlight its quality and diversity.
3.1 Empirical Analysis
Existing communication studies [15, 16] suggest that fake news videos can be classified from multiple dimensions, such as production motives, technical means, and content nature. We analyze both public fake news video datasets and real-world collected samples. Fake news often misleads through edited and spliced footage or emotionally manipulative, contradictory content. Since news videos typically contain three modalities—title, video, and audio—we examine the semantic differences among them. When one modality conflicts with the others or defies common sense, the video is more likely to be fake. Based on modality-level inconsistency, we categorize fake news videos into four types:
Figure 1: Example of Contextual Dishonesty: The video depicts children in Gaza longing for peace, while the title claims they anticipate continued conflict.
3.1.1 Contextual Dishonesty
Titles, as the video’s summary, significantly shape audience perception. Malicious title manipulation causes semantic dissonance by omitting context or using emotionally charged language. Common tactics include subject-object substitution, exaggeration, and removal of qualifiers—misleading readers who engage only with headlines.
Figure 2: Example of Cherry-picked Editing: A fabricated claim that Musk is buying GM, stitched together from unrelated clips.
3.1.2 Cherry-picked Editing
Video content is often altered by omitting key scenes or splicing unrelated footage to construct a false narrative. Tactics include removing moderate remarks, excluding peaceful scenes, or combining unrelated events, thereby amplifying bias and misleading viewers.
Figure 3: Example of Synthetic Voiceover: A party video overlaid with unrelated political commentary.
3.1.3 Synthetic Voiceover
These videos pair AI-generated or cloned speech with unrelated visuals. For example, a fake announcement might play over footage of a harmless event. The mismatch exploits viewers’ assumptions of audiovisual coherence, often bypassing detection through professional editing.
Figure 4: An example of Contrived Absurdity fake news video. The video shows a gigantic moon, and the title says it will disappear in 30 seconds, which clearly defies common sense.
3.1.4 Contrived Absurdity
Such videos use consistent multimodal content to present absurd claims (e.g., “alien bodies found” or “5-ton dog”), often generated by AI or CGI. Despite surface coherence, the narratives defy logic, aiming to attract attention through curiosity and shock.
3.2 Data Collection
At present, the existing fake news video detection datasets are collected from the short video platforms Douyin and Tiktok. These videos are mainly published by individual users with a small number of followers. Moreover, the clips of fake news videos are poor and people can easily distinguish the authenticity from the fake ones, so this type of fake news videos is not too harmful. The news video released by the news media is well-produced and will receive more attention, and the news media may also publish fake news videos for political purposes or commercial interests, which is particularly harmful. To solve this problem, we choose to collect news videos posted by news media from Twitter and YouTube. Specifically, we use web crawler technology to collect videos on 12 topics including accidents, epidemics and politics published by 27 news media on Twitter and YouTube in the past five years. In order to minimize noise and improve the quality of the dataset, we manually check and remove the videos with blurred images and too short duration. A total of 2,393 news videos are retained. These news videos contain information in three modals: title, video clip and audio, and the semantic information expressed in all three modals is the same.
Table 1: Examples of prompts and titles that are before and after modification.
Prompt Original Title Modified Title
Title: [title]. Change the meaning of the title from the perspective of object. A Survey Among Year 11 Students in London Reveals a Majority Would Voluntarily Wear Masks. A Survey Among Year 11 Students in New York Reveals a Majority Would Voluntarily Wear Masks
Title: [title]. Change the meaning of the title from the perspective of action. Duke of Edinburgh involved in a road traffic accident, but he is not injured. Duke of Edinburgh injured in Road Traffic Incident.
Title: [title]. Regenerate a news headline for a completely different event under the same topic A recent hit-and-run incident on Glebe Road just south of Route 50 near Goodwill. Multi-Vehicle Pileup Shuts Down I-95 During Morning Rush Hour, Injuries Reported.
3.3 Data Augmentation
Since our news videos are collected from news media, most of the videos are real, so there is a problem of data imbalance. Therefore, inspired by Liu et al. [13], we adopt LLMs to generate fake news videos from original real news videos. According to the four categories of fake news videos in the empirical analysis, we classify the fake news videos in FakeTT. Then we use different methods to generate the corresponding four different categories of fake news videos, and make the proportion of the four types similar to that in FakeTT, resulting in CD:600, CE:450, SV:300, CA:150.
Contextual Dishonesty (CD) Since the semantic information of the title mismatches other modalities, we utilize LLMs to modify the original video titles, ensuring inconsistency with the semantic content of the video. Specifically, we employ ERNIE 4.0 to regenerate video titles with divergent semantics. We design prompts to instruct the LLMs to alter either the object, verb phrase, or generate different events under the same topic in the titles. Table 1 illustrates our prompts and examples of modified titles. Subsequently, we manually review and remove redundant information from the generated responses, refining poorly constructed instances. This process yielded CD-type fake news videos with semantically mismatched titles.
Cherry-picked Editing (CE) Semantic inconsistencies caused by modifications in the video modality primarily arise from deleting key frames of the original material and splicing different segments. To simulate this forgery process, we adopt a similar approach. For selecting critical segments in the video, we use TransNetV2  [21] to perform shot segmentation on the video clips. Subsequently, frames are extracted from each video segment at two-seconds intervals. The video’s title and all extracted frames from each segment are then mapped into a shared vector space using CLIP. The cosine similarity between each frame and the title is calculated, and the average similarity across all frames in a segment is assigned as the segment score. After comparing the scores of all segments, we remove the video segments with the top 1/3 of the segment score. Due to the loss of critical information, these videos exhibit logical incoherence and modality mismatch, thereby being classified as CE-type fake news videos.
Synthetic Voiceover (SV) Semantic contradictions in audio modals are artificially created by generating audio content that diverges from the video’s original topic. To achieve this, we leverage LLMs to synthesize textually unrelated narratives under the same topic. Specifically, the video’s title and image captions of key frames are fed into ERNIE with structured prompts to generate topic-coherent but semantically irrelevant news articles. These generated texts are then converted into speech using text-to-speech tools VITS to replace the original audio. The resulting videos exhibit cross-modal contradictions between audio narration and visual content, classified as SV-type fake news videos.
Contrived Absurdity (CA) This type of fake news video maintains semantic consistency across modalities but violates common sense through illogical or absurd claims. Adopting a method similar to the CD-type processing, we utilize ERNIE 4.0 to exaggerate the video title in an illogical way while preserving their original meaning. The prompts instruct the LLMs to amplify specific elements (e.g., inflating quantities or distorting causality) without altering core entities or events. In this way, we get CA-type fake news videos with the same semantic information but contrary to common sense.
Following data augmentation, we end up with a final dataset comprising five categories of 2393 videos.
Table 2: Comparison of datasets for MFND. DY: Douyin; TT: Tiktok; YT: YouTube; TW: Twitter. The tokens number of title and the video duration is the average of the dataset.
Name Language Source Instances (Fake/Real) Title (Tokens) Duration Publisher
FakeSV Chinese DY 1827/1827 22.4 38.5s Users
FakeTT English TT 1172/819 21.8 48.1s Users
FMNV English YT&TW 1500/893 21.9 73.8s Media
3.4 Data Analysis
The FMNV dataset comprises a total of 2,393 news videos, including 893 videos categorized as real and 1,500 videos categorized as fake. The fake category can be further divided into four subtypes: CD: 600 videos with title modifications causing semantic inconsistencies; CE: 450 videos where key video clips matching the title were deleted; SV: 300 videos with audio replaced by reports of different events under the same theme; CA: 150 videos featuring titles with synonymous but exaggerated narration.
When compared to other MFND video datasets, FMNV offers notable advantages. In our LLM text generation process for data augmentation, we use a word substitution strategy to replace some words from the original text to change the original meaning. This ensures that the length of both positive and negative texts remains comparable, thereby preventing significant length disparities. With an average video duration of 73.8 seconds, FMNV boasts considerably longer videos than other datasets, and nearly half of the videos are long videos longer than 80 seconds, allowing for a richer information content per video. A comparative analysis of MFND datasets is presented in Table 2.
4 Method
Fig. 5 illustrates the overall architecture of the proposed FMNVD. Our baseline model consists of two core components: feature extraction and feature aggregation. The feature extraction module utilizes BERT to process textual features from video titles and audio transcripts, while employing a dual-stream architecture with Faster R-CNN and CLIP to capture both local object features and global visual representations from video frames. The feature aggregation module integrates these multimodal features through co-attention mechanisms that model cross-modal correlations, ultimately fusing the enhanced representations for fake news video classification.
4.1 Feature Extraction
4.1.1 Title
In the title feature extraction module, the video title text is first tokenized using the BERT-base WordPiece tokenizer with added [CLS] and [SEP] tokens. The token sequence is then fed into a 12-layer Transformer encoder with positional embeddings. We extract the final hidden states of all tokens from the last Transformer layer, where the [CLS] token’s 768-dimensional embedding serves as the global semantic representation. To handle variable-length inputs, sequences are padded/truncated to a fixed length $l_{t}=32$. The resulting text features $F_{t}=[w_{1},…,w_{l_{t}}]$ capture contextualized word-level semantics for downstream fusion.
Figure 5: Overview of proposed FMNVD.
4.1.2 Video
We employ temporal sampling to extract key frames from video sequences. Specifically, given a video $V$ with duration $T$, we uniformly sample $N$ frames at intervals of $\Delta t$ seconds. Each frame $I_{t}$ is resized to $H\times W$ resolution with RGB normalization.
Our dual-stream architecture synergistically combines local object details and global contextual patterns. For the local object features, Faster R-CNN with ResNet-101-FPN backbone processes each frame to generate region proposals through Region Proposal Network (RPN). We retain the top-$K$ proposals with confidence scores exceeding $\tau=0.5$, with $K=50$. followed by RoI pooling to extract fixed-size features $\{o_{i}\}_{i=1}^{K}\in{R}^{256}$. For global context features, CLIP (ViT-B/32) encoder partitions each frame into a $3\times 3$ grid of patches, yielding 9 patch embeddings $\{p_{j}\}_{j=1}^{9}\in{R}^{512}$ through linear projection.
Before fusion, we perform cross-modal feature alignment: 1) Project local features to CLIP’s semantic space via learnable matrix $W_{a}\in{R}^{512\times 256}$, 2) Apply element-wise L2-normalization to both modalities:
$\hat{o}_{i}=\frac{W_{a}o_{i}}{\|W_{a}o_{i}\|_{2}},\quad\hat{p}_{j}=\frac{p_{j}}{\|p_{j}\|_{2}}$ (1)
The final fused representation concatenates aligned features followed by dimension compression:
$F_{f}=\phi\left(\left[\hat{o}_{1},\dots,\hat{o}_{K}\oplus\hat{p}_{1},\dots,\hat{p}_{9}\right]\right)\in{R}^{d}$ (2)
where $\oplus$ denotes concatenation, $\phi:{R}^{(K+9)\times 512}\to{R}^{d}$ is a trainable projection layer (implemented as two-layer MLP with ReLU), and $d$ is the unified frames feature dimension.
To aggregate frame-wise features into a unified video representation, we employ a two-stage spatial-temporal fusion strategy. First, all frame features are concatenated along the temporal dimension to preserve raw forensic patterns. Then, a learnable compression module projects the high-dimensional features into a compact embedding space.
The video representation is formally obtained through:
$F_{v}=\psi\left(\bigoplus_{t=1}^{N}F_{f}^{(t)}\right)$ (3)
where $\bigoplus$ denotes full concatenation of $N$ frame features, and $\psi:{R}^{N\times 512}\to{R}^{d}$ represents the dimension reduction network implemented as:
$\psi(X)=W_{2}(\text{ReLU}(W_{1}X+b_{1}))+b_{2}$ (4)
with $W_{1}\in{R}^{k\times 512N}$, $W_{2}\in{R}^{d\times k}$ forming a bottleneck structure. Layer normalization is applied before each linear transformation to stabilize training.
4.1.3 Audio
We leverage semantic information from audio by first transcribing the audio stream into text using Whisper. The transcribed text is tokenized via BERT’s WordPiece tokenizer following the same pipeline as the title module, starting with a [CLS] token and ending with [SEP]. The sequence is padded/truncated to a fixed length $l_{a}=64$. Processed by the shared BERT encoder, the tokens are transformed into contextualized embeddings. We extract the final hidden states of the first 64 tokens as audio-derived text features $F_{a}=[w_{1},…,w_{l_{a}}]$, where parameter sharing with the title encoder ensures cross-modal representation alignment.
Figure 6: The main structure of Co-Attention.
4.2 Feature Aggregation
To effectively integrate information across modalities, we adopt a hierarchical co-attention mechanism inspired by Qi et al. [3], the main structure of Co-Attention is shown in figure 6. which enables deep interaction between textual, auditory, and visual signals. This design helps the model capture fine-grained cross-modal alignments and contextual dependencies.
We first apply co-attention between the title features $F_{t}\in{R}^{L_{t}\times d}$ and the audio transcription features $F_{a}\in{R}^{L_{a}\times d}$. Both modalities are projected into a shared space using linear layers. Then, each modality attends to the other by computing multi-head attention in both directions:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (5)
Specifically, $F_{t}$ acts as query and attends to $F_{a}$ (as key and value), and vice versa. The outputs from both directions are integrated with residual connections and layer normalization:
$F_{t}^{\prime}$ $=\text{LayerNorm}(F_{t}+\text{Attn}_{a\rightarrow t})$ (6)
$F_{a}^{\prime}$ $=\text{LayerNorm}(F_{a}+\text{Attn}_{t\rightarrow a})$ (7)
After audio-text fusion, we use the enhanced title representation $F_{t}^{\prime}$ to interact with visual frame features $F_{v}\in{R}^{L_{v}\times d}$ using the same co-attention procedure. This produces updated features $F_{t}^{\prime\prime}$ and $F_{v}^{\prime}$:
$F_{t}^{\prime\prime},F_{v}^{\prime}=\text{CoAtt}(F_{t}^{\prime},F_{v})$ (8)
The final fused representation is obtained by concatenating the three enhanced modalities:
$F=\text{Concat}(F_{t}^{\prime\prime},F_{a}^{\prime},F_{v}^{\prime})\in{R}^{3d}$ (9)
This fused vector is linearly projected and passed into a Transformer encoder with $L$ layers to capture higher-order interactions, followed by a multi-layer perceptron (MLP) for final classification:
$\hat{y}=\text{MLP}(\text{Transformer}(FW^{\top}))$ (10)
By gradually enhancing the title features through audio and visual interactions, our model captures richer multi-modal semantics and improves its ability to detect subtle inconsistencies indicative of fake news.
5 Experiments
In this section, we conduct experiments to evaluate the performance of FMNVD and other fake news detection methods on our newly proposed FMNV dataset.
5.1 Baseline
5.1.1 Pre-trained Models
1) BERT+ResNet We adopt a dual-stream architecture where the text modality is encoded using BERT [17], a transformer-based language model, to capture contextual semantics. For visual content, ResNet-50 [18] is utilized to extract spatial features from video frames or images. The text and visual features are concatenated or fused via cross-modal transformers to predict authenticity. 2)CLIP [19], a vision-language model pre-trained on 400 million image-text pairs. We fine-tune CLIP’s text and image encoders on the FMNV and compute similarity scores between video frames and accompanying text.
5.1.2 Neural Network-based Models
1) SV-FEND [3] The model leverages cross-modal attention mechanisms to enhance feature representation by capturing correlations between different modalities and integrates social context information to improve detection accuracy. 2) FakingRecipe [14] This model detects fake news in short videos by analyzing material selection and editing behaviors. It captures sentiment and semantic clues from audio, text, and visuals during material selection, and spatial-temporal editing patterns during video editing, enhancing detection accuracy.
5.2 Experimental Settings
The models are optimized using the Adam optimizer [20] with an initial learning rate of $1\times 10^{-3}$. To ensure a rigorous evaluation, we measure four classification metrics: Accuracy, F1-score, Precision, and Recall. All experiments are conducted on an NVIDIA GeForce RTX 2080 Ti GPU with a batch size of 128 for 30 epochs. The cross-entropy loss function is minimized during training, defined as:
$\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(p_{i,c})$ (11)
where $N$ is the batch size, $C$ is the number of classes, $y_{i,c}$ denotes the ground-truth label, and $p_{i,c}$ represents the predicted probability for class $c$.
5.3 Experiments Results
Table 3: Performance of different methods on FMNV.
Methods Test Data Accuracy F1 Precision Recall
BERT+ResNet CD CE VS CA ALL 57.67 78.01 78.97 73.40 64.17 52.41 76.72 75.26 66.57 63.80 53.95 75.39 74.69 65.40 64.83 53.17 79.93 79.81 69.80 65.78
CLIP CD CE VS CA ALL 58.88 79.62 80.08 76.40 69.35 53.93 77.83 77.35 65.62 69.21 55.29 77.25 76.10 64.40 69.82 54.35 80.41 83.99 69.21 69.91
SV-FEND CD CE VS CA ALL 60.00 83.70 82.50 76.19 72.92 55.44 83.13 80.43 65.73 72.26 56.94 83.13 79.41 64.45 72.26 55.83 87.22 88.33 75.00 73.67
FakingRecipe CD CE VS CA ALL 60.67 88.15 88.33 82.85 73.75 52.88 87.39 86.32 68.42 72.70 57.27 86.44 84.09 66.98 72.16 54.72 90.00 92.22 70.56 72.77
FMNVD CD CE VS CA ALL 64.00 88.15 86.67 80.95 74.17 58.64 87.49 84.37 69.12 72.68 62.06 86.63 82.30 66.88 72.52 59.17 90.56 90.00 75.00 72.89
To validate the effectiveness of our proposed baseline and other FNVD methods on the FMNV dataset, we designed the following experiments: We categorize the fake samples in the test set into four groups - CD, CE, SV, and CA - corresponding to four distinct data augmentation methods. We evaluate several models on each categorical test subset to analyze their detection capabilities for different types of forged news videos. Additional evaluation was conduct on the entire test set to verify the overall effectiveness of our proposed method. The experimental results shown in Table 3 lead us to the following conclusions: 1) All three neural network-based methods achieved accuracies above 0.72 on the FMNV dataset, with FMNVD demonstrating superior performance among the five baseline models by attaining 0.7417, which confirms both the generalization capability of FMNV and the competitive advantage of FMNVD. 2) The evaluated methods exhibited excellent detection performance for fake news videos containing Cherry-picked Editing and Synthetic Voiceover manipulations. However, they show relatively poor detection accuracy for Contextual Dishonesty-type videos. This limitation primarily stems from the fact that Contextual Dishonesty samples are generated through subtle phrase-level modifications in video titles, which creates semantically nuanced variations that current models struggle to detect effectively. This presents a significant challenge for future methodological improvements.
Table 4: Ablation study on FMNVD. (a) Modality removal. (b) Hyperparameter sensitivity.
(a) Modality Ablation
Data Accuracy F1 Precision Recall
All 74.17 72.68 72.52 72.89
w/o Title 65.42 65.37 68.16 68.78
w/o Frames 71.25 70.47 70.43 71.67
w/o Audio 63.75 58.83 60.22 58.78
(b) Hyperparameter Sensitivity
Setting Accuracy F1 Precision Recall
Patch size = 16 73.84 72.23 72.00 72.50
Seq. len. = 32 74.17 71.91 72.46 71.56
Seq. len. = 64 73.75 72.96 72.80 74.11
$\tau=0.3$ 74.17 71.12 72.82 70.44
$\tau=0.7$ 72.50 70.67 70.67 70.67
5.4 Ablation Study
To assess both the contribution of each modality and the robustness of our framework to key hyperparameters, we perform two sets of ablations on FMNVD:
(a) Modality Ablation We individually remove each modality (title, frames, audio) and its associated co-attention layers, and measure the resulting drop in performance. As shown in the top block of Table4, all three modalities are vital, with audio having the largest impact, followed by title and then visual frames. This phenomenon primarily stems from our implementation strategy where audio signals are first converted into text through Whisper for processing. The transcribed audio text inherently contains more linguistic content than video titles, and such extended textual information proves more impactful than shorter textual inputs (titles) and visual features in our framework.
(b) Hyperparameter Sensitivity To verify the stability of our detection framework, we further vary three core hyperparameters:
  • CLIP patch size (e.g. 16)
  • Maximum sequence length for text/audio features (e.g. 32 vs. 64 tokens)
  • Confidence threshold $\tau$ for filtering low-confidence predictions (e.g. 0.3 vs. 0.7)
The bottom block of Table 4 provides the performance under each setting, demonstrating that our model maintains stable accuracy and F1 across a reasonable range of each hyperparameter.
6 Conclusion
In this paper, we conduct an empirical analysis for fake news video categorization and construct a novel fake news video dataset (FMNV) using large model-based data augmentation. We provide a comprehensive description of FMNV’s creation process with in-depth investigations into data entity attributes and distribution characteristics. To assess the dataset’s validity, we introduce a new baseline model FMNVD and perform numerical experiments to evaluate various baseline performances, thereby demonstrating FMNV’s generalization capability. Future work will explore additional promising approaches to enhance the dataset’s predictive power, with the expectation that FMNV will pave the way for advanced developments in MFND research.